A local bus traveling at a rate of 50 kilometers per hour was passed by an express bus traveling at a rate of 75 kilometers per hour. The express bus left the terminal one hour later than the local bus. How far had the buses traveled?
Answer by checkley71(8403)

50X=75(X-1)
50X=75X-75
50X-75X=-75
-25X=-75
X=-75/-25
X=3 HOURS AFTER THE FIRST BUS LEFT THE FASTER BUS OVERTOOK IT.
PROOF:
50*3=75(3-1)
150=75*2
150=150
---------------------------------------------------------------------
Thank you so much for your help with this. Your reply was very quickly received. I have a follow up question. What is X standing for? If I read it correctly, we first have to solve for how much time it took for the express bus to overtake the public bus, and then once we have the time, we can determine the distance? Am I understanding correctly?
------------------------------------------
ANSWER YES!!!!!!